# Differentiable Physics-based System Identification for Robotic Manipulation of Elastoplastic Materials

### v2

@@ Abstract
Robotic manipulation of volumetric elastoplastic deformable materials, from foods such as dough to construction materials like clay, is in its infancy, largely due to the difficulty of modelling and perception in a high-dimensional space. Simulating the dynamics of such materials is computationally expensive. It tends to suffer from inaccurately estimated physics parameters of the materials and the environment, impeding high-precision manipulation. Estimating such parameters from raw point clouds captured by optical cameras suffers further from heavy occlusions. To address this challenge, this work introduces a novel Differentiable Physics-based System Identification (DPSI) framework that enables a robot arm to infer the physics parameters of elastoplastic materials and the environment using simple manipulation motions and incomplete 3D point clouds, aligning the simulation with the real world. Extensive experiments show that with only a single real-world interaction, the estimated parameters, Young’s modulus, Poisson’s ratio, yield stress and friction coefficients, can accurately simulate visually and physically realistic deformation behaviours induced by unseen and long-horizon manipulation motions. Additionally, the DPSI framework inherently provides physically intuitive interpretations for the parameters in contrast to black-box approaches such as deep neural networks.
@@ 3.3 Differentiable dynamics modelling
In particular, for each numerical computation step, DiffTaiChi flattens its computation branches (e.g., boundary and collisions) and replaces mutable local variables with extra local storage variables, producing straight-line codes without mutable variables. It then uses standard source code transform to generate the derivative function based on the adjoint method. To compute gradients with a loss function, DiffTaiChi records the order of computation kernels and the scalar variables in the forward simulation direction, and then it computes the gradients for the concerned variables by evaluating the derivative functions in the reversed simulation direction (Hu et al., 2020). We build a simulator based on this programming language to allow automatic gradient computation for the physics parameters. The codes for the simulator will be open-sourced upon paper acceptance.
@@ 3.4 Real-to-sim object reconstruction
Real-world platform: To collect real-world data, we set up a deformable object manipulation system where a Kuka IIWA LBR 14 industrial arm (Kuka, 2024) is equipped with a Zivid One+ medium camera (Zivid, 2024) for perceiving the object, and one of the three end-effectors, namely a rectangular cuboid, a cylinder roller and a bullet-shaped object for collecting interaction experiences with different contact geometries, as shown at the bottom-left of Figure 1. We use plasticine (non-hardening modelling clay) as the main material for our experiments.
@@ 4.3 In-distribution performances
Secondly, by examining the details of the heightmaps more carefully, it shows that optimising the CD losses tends to produce heightmaps with darker colours (which represent greater heights of the object) while optimising the EMD losses tends to produce heightmaps where the objects look slightly bigger and wider. It indicates that optimising the CD losses tends to produce particles that match the height of the ground truth shapes, i.e., in the $z$ direction, while optimising the EMD losses tends to match the $x$ and $y$ directions. This is especially true at level-1 contact complexity, where the heightmaps in the first two rows for each object configuration in Subfigures 6(a) and 6(b) are generally darker than those in the last two rows. At level 2, the PRT CD loss sometimes changes its focus onto the $x$ and $y$ directions, producing lighter-colour heightmaps (the second rows of the heightmaps associated with the rectangle and round end-effectors in Subfigures 6(c) and 6(d)).
Thirdly, though the particles and heightmaps are very similar, they are produced by quite different parameter combinations. For example, Table 4 reveals a positive correlation between the yield stress ($\boldsymbol{\sigma}_{y}$) and material density ($\rho$). One can see that, although the simulated results are similar, a larger $\boldsymbol{\sigma}_{y}$ is always accompanied by a larger $\rho$, and vice versa. This is realistic because when the object becomes heavier (greater $\rho$), it needs to be more difficult to yield (deform plastically); otherwise, it will collapse due to gravity. If the object is light (lower $\rho$), a smaller yield stress value allows it to exhibit the returning effect after compression.
In addition, from the results at contact level 2 in Subfigures 6(c) and 6(d), one can also observe a correlation between the friction coefficients and the material density – the three key parameters that determine how much the object will be moved in the shifting motion direction. The parameters in Table 4 and their corresponding visualisations show that,
@@
Fourthly, the visualisations at both levels also reveal limitations and potential improvement directions of the physics model. By comparing with the ground truths, the simulated contact areas of the objects always deform more sharply with insufficient elastic returning. The real objects, on the other hand, tend to respond more elastically after being plastically deformed, hence the higher and smoother surfaces shown in the ground truths. This is more obvious for contact areas with sharp edges such as the four edges of the rectangle end-effector or the sharp sides of the cylinder end-effector. In addition, at level-2 cases, with the effectors shifting horizontally, some simulated particles are unrealistically displaced and stay floating. In the real world, the displaced parts would fall because of gravity. What’s more, the shifting motion in the real world causes the whole object to tilt in the moving direction, while in the simulation the contact impact tends to remain in a much smaller region around the contact area. These inconsistencies can be caused not only by inaccurate modelling, but also by various computation approximations, such as time integration method, simulation step size, contact handling, etc.
@@
  • There are many local minima and possible parameter solutions, but they produce visually and physically similar manipulation results.
@@ 4.4 Generalisation
Secondly, for the unseen and more complex motions, the found physics parameters still retain their physical meanings and interpretability. Evidence supporting this can be found by observing the differences between the heightmaps in Figure 8 produced by different parameters.
  • The first example is that, when the yield stress ($\boldsymbol{\sigma}_{y}$) is smaller, the resultant heightmaps of the unseen motions tend to be lighter, meaning the objects are more compressed down.
Thirdly, the found solutions are not globally optimal. Despite the visual similarity, Table 5 shows that the best generalisation results are not always achieved by the cases with the best in-distribution validation performances, as the underlined and bolded texts rarely coincide. In addition, though the hand-picked parameters do achieve the best results in two out of six object simulations, they are not the best fit for every scenario. This again suggests the importance of re-fitting the parameters on the fly for every new scenario as there are hardly two objects that are identical especially for elastoplastic materials. Qualitatively, from the heightmaps and trajectories in Figures 8 and 7, one can also recognise certain visual discrepancies between the real objects and their simulation counterparts, indicating a certain degree of mismatch between the found parameters and the actual solutions. This could be improved by further diversifying the data (and the motions that produce them) used for optimisation. Other directions for improvement may include better observational noise removal or better physics model selection/approximation.
Finally, the modelling limitations observed in the in-distribution visualisation still exist in the unseen motion visualisations (Figs. 7 and 8). The insufficient elastic returning at contact areas especially with sharp manipulator edges can be seen from all visualisations where the primary contact areas of the object rarely return to heights close to the real object. Floating displaced particles can be observed from the poking-180-rotating motion visualisations where many particles are carried away during rotations. The insufficient spreading of the influence of contacts can be recognised from the first object configuration of the flattening motion visualisations where the object is not tilted enough in the second half of the trajectory.
@@ ACKNOWLEDGMENT
This work was supported by the Engineering and Physical Sciences Research Council (grant No. EP/X018962/1).

### v3

@@ Abstract
Robotic manipulation of volumetric elastoplastic deformable materials, from foods such as dough to construction materials like clay, is in its infancy, largely due to the difficulty of modelling and perception in a high-dimensional space. Simulating the dynamics of such materials is computationally expensive. It tends to suffer from inaccurately estimated physics parameters of the materials and the environment, impeding high-precision manipulation. Estimating such parameters from raw point clouds captured by optical cameras suffers further from heavy occlusions. To address this challenge, this work introduces a novel Differentiable Physics-based System Identification (DPSI) framework that enables a robot arm to infer the physics parameters of elastoplastic materials and the environment using simple manipulation motions and incomplete 3D point clouds, aligning the simulation with the real world. Extensive experiments show that with only a single real-world interaction, the estimated parameters, Young’s modulus, Poisson’s ratio, yield stress and friction coefficients, can accurately simulate visually and physically realistic deformation behaviours induced by unseen and long-horizon manipulation motions. Additionally, the DPSI framework inherently provides physically intuitive interpretations for the parameters in contrast to black-box approaches such as deep neural networks. The project is fully open-sourced via GitHub.
@@ 3.3 Differentiable dynamics modelling
In particular, for each numerical computation step, DiffTaiChi flattens its computation branches (e.g., boundary and collisions) and replaces mutable local variables with extra local storage variables, producing straight-line codes without mutable variables. It then uses standard source code transform to generate the derivative function based on the adjoint method. To compute gradients with a loss function, DiffTaiChi records the order of computation kernels and the scalar variables in the forward simulation direction, and then it computes the gradients for the concerned variables by evaluating the derivative functions in the reversed simulation direction (Hu et al., 2020). We build a simulator based on this programming language to allow automatic gradient computation for the physics parameters.
@@ 3.4 Real-to-sim object reconstruction
Real-world platform: To collect real-world data, we set up a deformable object manipulation system where a Kuka IIWA LBR 14 industrial arm (Kuka, 2024) is equipped with a Zivid One+ medium camera (Zivid, 2024) for perceiving the object, and one of the three end-effectors, namely a rectangular cuboid, a cylinder roller and a bullet-shaped object for collecting interaction experiences with different contact geometries, as shown at the bottom-left of Figure 1. We use plasticine (non-hardening modelling clay) as the main material for our experiments and cloud slime and soil for generalisation experiments. The clay and cloud slime are purchased from Amazon [^1] Modelling Clay. Cloud Slime Kit.
@@ 4.3 In-distribution performances
Secondly, by examining the details of the heightmaps more carefully, it shows that optimising the CD losses tends to produce heightmaps with darker colours (which represent greater heights of the object) while optimising the EMD losses tends to produce heightmaps where the objects look slightly bigger and wider. This is especially true at level-1 contact complexity, where the heightmaps in the first two rows for each object configuration in Subfigures 8(a) and 8(b) are generally darker than those in the last two rows (see figure 6 for enlarged examples). At level 2, the PRT CD loss sometimes changes its focus on the $x$ and $y$ directions, producing lighter-colour heightmaps (see the second rows of the heightmaps associated with the rectangle and round end-effectors in Subfigures 8(c) and 8(d)). This observation indicates that optimising the CD losses tends to produce particles that match the height of the ground truth shapes, i.e., in the $z$ direction, while optimising the EMD losses tends to match the $x$ and $y$ directions.
Figure 6: Selected examples from Figure 8(b) to show that the CD losses produce darker-coloured heightmaps while the EMD losses produce lighter-coloured and wider heightmaps, which can be intuitively explained by the optimised yield stress values and their potential difference in spatial focus.
Figure 7: Selected examples from Figure 8(d) to show the physically realistic associations between object movements and three physics parameters, demonstrating the physical plausibility of different physics parameter values.
@@
Thirdly, though the particles and heightmaps are very similar, they are produced by quite different parameter combinations. From the results at contact level 2 in Subfigures 8(c) and 8(d), one can observe a correlation between the friction coefficients and the material density – the three key parameters that determine how much the object will be moved in the shifting motion direction. The parameters in Table 3 and their corresponding visualisations exhibit the following physically plausible relationships (see Figure 7 for a few examples),
@@
Fourthly, the visualisations in Figure 8 also reveal limitations and potential improvement directions of the physics model. By comparing with the ground truths, the simulated contact areas of the objects always deform more sharply with insufficient elastic returning. The real objects, on the other hand, tend to respond more elastically after being plastically deformed, hence the higher and smoother surfaces shown in the ground truths. This is more obvious for contact areas with sharp edges, such as the four edges of the rectangle end-effector or the sharp sides of the cylinder end-effector. In addition, at level-2 cases, with the effectors shifting horizontally, some simulated particles are unrealistically displaced and stay floating. In the real world, the displaced parts would fall because of gravity. What’s more, the shifting motion in the real world causes the whole object to tilt in the moving direction, while in the simulation the contact impact tends to remain in a much smaller region around the contact area. These inconsistencies can be caused not only by inaccurate modelling, but also by various computation approximations, such as time integration method, simulation step size, contact handling, etc.
@@
  • There are many local minima and possible parameter solutions, but they produce visually and physically similar manipulation results with physically intuitive interpretability.
@@ 4.4 Generalisation
Secondly, the found physics parameters retain their physical meanings and interpretability for the unseen and more complex motions. Evidence supporting this can be found in Figure 12 produced by different parameters (see Figure 10 for selected examples).
  • The first example is that when the yield stress ($\boldsymbol{\sigma}_{y}$) is smaller, the resultant heightmaps of the unseen motions tend to be lighter, meaning the objects are more compressed down.
Figure 10: Selected examples from Figure 12 to show the physically realistic associations between the manipulation results and the found parameter values. Comparisons to the ground truths reveal the modelling limitations (insufficient elastic returning).
Figure 11: Selected frames from two trajectories in Figure 9 to show modelling limitations (floating particles and insufficient spreading of contacts). The areas around the green cycles show that the second part of the flattening motion does not tilt the object up, and there are floating particles in the poking-180-rotating motion.
@@
Thirdly, the found solutions are not globally optimal. Despite the visual similarity, Table 4 shows that the best generalisation results are not always achieved by the cases with the best in-distribution validation performances, as the underlined and bolded texts rarely coincide. In addition, though the hand-picked parameters do achieve the best results in two out of six object simulations, they are not the best fit for every scenario. This again suggests the importance of re-fitting the parameters on the fly for every new scenario as there are hardly two objects that are identical especially for elastoplastic materials. Qualitatively, from the heightmaps and trajectories in Figures 12 and 9, one can also recognise certain visual discrepancies between the real objects and their simulation counterparts, indicating a certain degree of mismatch between the found parameters and the actual solutions. This could be improved by further diversifying the data (and the motions that produce them) used for optimisation. Other directions for improvement may include better observational noise removal or better physics model selection/approximation.
Finally, the modelling limitations observed in the in-distribution visualisation still exist in the unseen motion visualisations (Figures. 9 and 12). The insufficient elastic returning at contact areas, especially with sharp manipulator edges, can be seen from all visualisations where the primary contact areas of the object rarely return to heights close to the real object (compare also the enlarged heightmaps in Figure 10). Floating displaced particles can be observed from the poking-180-rotating motion visualisations, where many particles are carried away during rotations. The insufficient spreading of the influence of contacts can be recognised from the first object configuration of the flattening motion visualisations where the object is not tilted enough in the second half of the trajectory. See Figure 11 for enlarged examples.
@@ ACKNOWLEDGMENT
This work was supported by the Engineering and Physical Sciences Research Council (grant No. EP/X018962/1). The authors would like to thank Samuel Moeller (Engineer support staff at the School of Engineering, Cardiff University) and Dr Prasad Rayamane for their help with 3D printing The end-effectors.